During an isothermal compression, water pressure increases from 33 atm to 66 atm. The water temperature remains constant at 20 C. If the isothermal compressibility is 5.8x10^-5 atm -1, the increase im density is?
Answer #1

W -5 em atml compressio =2=5,8X10 latm = 101.32 kPa dy Be dp dp = 66-33= 33 atm -dr 5.8x105 X(33) -5 191.4 XIO P.-dV 띠 م ہے -